# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v2

@@ 5 Visualizing results by LiePose diffusion [19]
We show some of LiePose [19] distribution estimates against our ground truth in Figure 8. Similarly to SpyroPose [9], LiePose [19] is able to retrieve the single mode of Instance 1, but gets better results when estimating continuous distributions.
@@
Figures 9, 10, 11 and 12 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries. SpyroPose [9] rotations tends to be more precise than LiePose [19], but misses some of the modes. LiePose [19] tends to estimate continuous symmetries when the image produces discrete ones. These images are taken from a video compilation of SpyroPose [9] and LiePose [19] results also provided as supplementary material (BOP_D_id239_supp_distribution_comparison_SpyroPose_LiePose.mp4). Scenes with single instance of objects have been chosen, to facilitate visualization. We invite the reader to stop on some frames and check the differences in the estimates. Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
@@
Figure 8: Illustration of LiePose [19] results.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.
@@
Figure 9: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 27 (four rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods are able to retrieve the four rotation modes of the object. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
@@
Figure 10: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 15 (continuous rotation). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods fail to generate the target continuous rotation, although SpyroPose produces more correct rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
@@
Figure 11: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 1 (six rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of six rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
@@
Figure 12: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 3 (eight rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of eight rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
@@ 6 Using Our Per-Image Annotations for Other Pose Estimation Datasets
We give here an illustration of the interest to reprocess their symmetries patterns. We take ITODD’s small validation set, for which the ground truth is public. Figure 13 presents our result for the star object new ground truth, as well as one case of the current best performer (gpose2023) failing to align the holes. In the current version of BOP evaluation, this pose estimation is validated. With our annotations, it would be penalized.
Figure 13: Illustration of ITODD symmetries on the validation set (with public GT). For the star image (left, first row), BOP symmetries display 12 rotation modes (middle), whereas our method would keep only 2 rotation modes, which align the two holes (right) (size was set to one over the number of modes, hence the bigger modes on the right). We also show a pose estimate of GPose [48], ranked first at BOP 2023, for the star object overlayed on its image. We observe that the holes are not correctly aligned. MSPD and MSSD metrics validate this estimate, whereas MSPD and MSSD metrics with our new annotations would have penalized it.

### v3

@@ 6 Visualizing results by LiePose diffusion [19]
We show some of LiePose [19] distribution estimates against our ground truth in Figure 14 on T-LESS. Similarly to SpyroPose [9], LiePose [19] is able to retrieve the single mode of Instance 1, but gets better results when estimating continuous distributions.
@@
Figures 15, 16, 17 and 18 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries. SpyroPose [9] rotations tends to be more precise than LiePose [19], but misses some of the modes. LiePose [19] tends to estimate continuous symmetries when the image produces discrete ones. These images are taken from a video compilation of SpyroPose [9] and LiePose [19] results also provided as supplementary material (BOP_Distrib_id8513_supp_distribution_comparison_SpyroPose_LiePose.mp4). Scenes with single instance of objects have been chosen, to facilitate visualization. We invite the reader to stop on some frames and check the differences in the estimates. Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
@@
Figure 14: Illustration of LiePose [19] results on T-LESS.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.
@@
Figure 15: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 27 (four rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods are able to retrieve the four rotation modes of the object. The envelop in the rotation part represents our BOP-Distrib annotation. We provide much more examples in the accompanying video.
@@
Figure 16: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 15 (continuous rotation). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods fail to generate the target continuous rotation, although SpyroPose produces more correct rotations. The envelop in the rotation part represents our BOP-Distrib annotation. We provide much more examples in the accompanying video.
@@
Figure 17: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 1 (six rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of six rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-Distrib annotation. We provide much more examples in the accompanying video.
@@
Figure 18: Visualizing SpyroPose [9] (top row) and LiePose [19] (bottom row) distribution results for object 3 (eight rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of eight rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-Distrib annotation. We provide much more examples in the accompanying video.
@@ 7 Using Our Per-Image Annotations for Other Pose Estimation Datasets
We give here an illustration of the interest to reprocess their symmetries patterns. We take ITODD’s small validation set, for which the ground truth is public. Figure 19 presents our result for the star object new ground truth, as well as one case of the current best performer for Single Pose estimation (gpose2023 [49]) failing to align the holes. In the current version of BOP evaluation, this pose estimation is validated. With our annotations, it would be penalized.
Figure 19: Illustration of ITODD symmetries on the validation set (with public GT). For the star image (left, first row), BOP symmetries display 12 rotation modes (middle), whereas our annotation method keeps only 2 rotation modes (right), which align the two holes (size was set to one over the number of modes, hence the bigger modes on the right). We also show (second row left) a pose estimate of GPose [49], ranked first at BOP 2023, for the star object overlayed on its image. We observe that the holes are not correctly aligned. MSPD and MSSD metrics validate this estimate, whereas MSPD and MSSD metrics with our new annotations would have penalized it.